OpenStudy (anonymous):

Can anyone help on solving equations for radical expressions? x + (the square root of x+6) = 9-x

OpenStudy (anonymous):

I understand that I must isolate the radical, meaning I should subtract the x from the right making the problem read the square root of x+6 = 9-2x Is this correct? From there I know to square both sides to rid the square root. So now the problem reads x+6 = 9-2x. Is this correct? opps! I mean x+6 = (9-2x)squared

OpenStudy (anonymous):

youre correct. if you ever need help with math go to Khanacademy.org. It is very helpful! Good luck and God bless!!!!!!!!!!!!!!!!!! :)